Documentation for poster.tex/poster.sty Posters and banners with Generic TEX

Version 0.91 January 30, 1993

Timothy Van Zandt tvz@Princeton.EDU

poster.tex/poster.sty contains the macro
    \poster{stuff }

for making posters and banners. stuff is processed in restricted horizontal mode (i.e., \hbox or “LR-mode”) and is then printed on as many sheets of paper as are needed. You can then construct the poster or banner by trimming and piecing together the sheets of paper.

You can also write

    \Poster stuff \endPoster

and LaTEX users can write

    \begin{Poster} stuff \end{Poster}

Here are a few details:

• Use a \vbox or LaTEX’s minipage or \parbox in stuff if you want to include vertical mode material.

• Don’t worry about margins, headers or footers; \poster ignores output routines entirely.

• It is up to you to make everything big. The easiest way is to set TEX’s \mag parameters (to an integer equal to 1000 times the magnification factor, or to \magstephalf, or to \magstepn, where n is 1,…,5). You may need to generate big bitmaps if using bitmapped fonts (e.g., TEX’s usual Metafont fonts), rather than scalable outline fonts (e.g., PostScript fonts). Using \magstepn minimizes the need for extra bitmaps.

• stuff can contain \catcode changes.

\poster and \Poster use the following parameters:

Parameter  Value  Default

paperwidth dim  8.5in
paperheight dim  11in
imagewidth dim  7.5in
imageheight dim  10in
landscape    true/false         false
hcenter      true/false         false
vcenter      true/false         false
crop         none/corners/full  corners
clip         none/pstricks      none

You can include parameter changes as a list of key=value pairs in an optional argument to \poster or \Poster, enclosed in square brackets. E.g.,

\poster[clip=pstricks,hcenter=true]{foo}

No extraneous spaces, please.

You can also redefine the parameters using \def or \LaTeX’s \renewcommand. For parameter foo, you should redefine \POSTERfoo. E.g., the next example is like the last one:

\def\POSTERclip{pstricks}
\def\POSTERhcenter{true}
\poster{foo}

This is mainly of interest when developing your own custom \poster com-

mand or poster.tex file.

Here are some comments on the parameters:

• Don’t adjust the page and image dimensions for your document’s magnification (and don’t use TEX’s true dimensions). poster.tex does this for you.

• Most printers cannot print right up to the edge of the paper. That is why the imagewidth and imageheight should be smaller than the paperwidth and paperheight. The default values are good for printing on 8.5in by 11in paper in portrait mode with one-half inch margins.

• Setting landscape to true is just a convenient way to switch the height and width parameters. You still have to take care of printing your document in landscape mode.

• Setting hcenter and vcenter to true causes the image to be centered horizontally and vertically, respectively, in the total number of pages that are printed. E.g., when vcenter is true, extra space is added to the top of the first row of pages and to the bottom of the last row of pages.

• The crop parameter controls the crop marks, to help you trim each page to size. When crop equals full, you get crop marks along the full length of all four sides.

• Setting clip to pstricks causes each page to be clipped to the size of the image (rather than having the image overlap in the margins on each page), but this only works if you have loaded the PSTricks package.

The file poster1.tex contains the following sample of a framed poster with a whole page of text:

\documentstyle[poster]{article}
\mag\magstep5  % Magnification of 1.2^5 (roughly 2.5)
% Use `true' dimensions below for magnified values.
\begin{document}
\begin{Poster}[vcenter=true,hcenter=true]
\setlength{\fboxsep}{.8truein}%
\setlength{\fboxrule}{.1truein}%
\fbox{\begin{minipage}{11.1truein}

stuff

\end{minipage}}%
\end{Poster}
\end{document}

The file poster2.tex contain the following sample of a banner in landscape mode. ptmr is meant to be the name of the Times-Roman PostScript font, if your TEX supports such a thing.

\font\bigroman=ptmr at 7.5in
\bigroman
\poster[vcenter=true,landscape=true]{Animals}